OpenStudy (anonymous):

Solve for x: 3 over x minus 2 over 3 equals 5 over 6 HELP!11

OpenStudy (shamil98):

3/x - 2/3 = 5/6 3/x = 5/6 + 4/6 3/x = 3/2 3 = 3/2 x 3(2/3) = x 6/3 = x 2 = x

OpenStudy (anonymous):